# NICER Spectral and Timing Analysis of 4U 1630–47 and its Heartbeat State

### v1

@@ 4.2 Relativistic thermal component
Figure 15: As in Fig. 14 but for fits with the spectral model tbfeo $\times$ thcomp $\otimes$ kerrd. Here, the inner disk radius ($R_{\mathrm{in}}$) is shown in the upper panel and the mass accretion rate is shown in the lower panel. $R_{\mathrm{in}}$ is in arbitrary unit due to a lack of robust constraint on the distance and mass of the source (Seifina et al., 2014).
@@ 5.1 Energy-resolved PDS
We further explore the timing properties of the heartbeat state, first computing the PDS for photons in different energy bands. We fit each with a Lorentzian model, which is commonly used to fit PDS including those with QPOs (Bachetti & Huppenkothen, 2022). For the PDS in the heartbeat state, the variability can be modeled with three Lorentzian functions. We set a zero-centered Lorentzian for the broad band noise, one peaked at the heartbeat characteristic frequency, and a third to fit the noise at higher frequencies.
@@ 5.2 Energy-resolved time lag and coherence
Analysis of the lags between the photons of different energies is a powerful tool to study the origin of variability in data (Wang et al., 2021; Bollemeijer et al., 2024; Ma et al., 2021), including heartbeats (Yang et al., 2022). In addition, coherence measures the degree of linear correlation between two time series (Bachetti & Huppenkothen, 2022), enabling us to assess the strength of the correlation between the oscillations of light curves of different energy bands. Here we choose the reference energy band to be 2–3 keV. We use stingray to calculate the frequency dependence of the lag and coherence of the chosen energy bands (3–4 keV, 4–7 keV, 7–12 keV) versus the reference energy band in the heartbeat states in 2021 and 2023, and the results are shown in Fig. 17. From the coherence, we can see that the light curves of different energy bands are relatively highly linearly correlated (coherence $\gtrsim 0.8$) near the characteristic frequency of the heartbeats annotated by the yellow region. A significant hard lag of $\sim 1~{}s$ between $4-12$ keV photons and 2–3 keV photons is clearly observed near the heartbeat frequency.
Figure 17: The energy dependence of lag (upper panels) and coherence (lower panels) of the heartbeat states in 2021 (left) and 2023 (right). The reference energy band is 2–3 keV, and the comparison bands are 3–4 keV, 4–7 keV, and 7–12 keV. The yellow region highlights the frequency range of the heartbeat oscillations (from the PDS fits). Near the characteristic frequency of the heartbeats annotated by the yellow region, a significant hard lag of $\sim 1~{}\mathrm{s}$ between 4–12 keV photons and 2–3 keV photons is clearly observed.
@@ 6 Discussion
6.1 The change of the inner disk radius in different states
@@
While at lower temperatures the disk emission can generally be well described by a multi-temperature disk blackbody spectrum, non-thermal effects such as the Compton scattering in the disk atmosphere may produce deviations from the multi-temperature disk blackbody spectrum, especially when the inner disk temperature is high. This is normally taken into account by introducing a color correction factor $f_{\mathrm{col}}$ (also called spectral hardening factor) (Shimura & Takahara, 1995; Kubota et al., 1998). Moreover, it has been found in previous studies that the color correction factor has strong dependence on the disk temperature (Kubota & Makishima, 2004; Davis et al., 2006), which can be approximately described by $f_{\mathrm{col}}\propto T_{\mathrm{in}}^{1/4}$. Based on this scaling relation and the equation $norm_{\mathrm{diskbb}}=(R_{\mathrm{in}}/f_{\mathrm{col}}^{2}D_{10})^{2}\cos{i}$ (Kubota et al., 1998), we overlay the constant $R_{\mathrm{in}}$ contour in Fig. 12, showing that $R_{\mathrm{in}}$ is essentially constant when the system is in the soft state. Similar results have also been found in Swift J1727–1613 (König et al. 2025, in preparation). We note a long-history of evidence from the thermal disk emission for soft states exhibiting a stable accretion-disk inner radius (e.g., detailed analysis of several decades of observations of LMC X–3; see Steiner et al. (2010), and references therein).
@@ 6.2 Heartbeat mechanism
Inner disk radiation instability is a popular candidate mechanism for producing the heartbeat oscillation (Lightman & Eardley, 1974; Done et al., 2007; Neilsen et al., 2012). In that scenario, the dominance of the radiation pressure in the inner disk region and the lack of balancing mechanism gives rise to a thermal-viscous instability cycle resulting in the heartbeat oscillations in the light curve (Done et al., 2007; Neilsen et al., 2011).
@@
To gain a better understanding of different physics processes behind this cycle behavior, it is useful to refer to the “S-curve” showing the relationship between the mass and temperature at a certain disk radius (e.g., see Fig. 7 in Done et al., 2007). The curve has three branches: a lower viscous slow rise (where the heating is balanced by radiative cooling and the temperature increases slowly with mass), an upper slow viscous fall (where the heating is balanced by advective cooling and the temperature decreases slowly with mass) and the middle thermal viscous cycle (where the accretion process switches quickly between the two stable branches) (Wang et al., 2024). This cycled oscillation causes a density wave to propogate along the disk and results in the fast changes in the mass accretion rate, inner disk radius, and temperature. The fast variability of disk parameters in our phase-resolved fitting results (Sect. 4) is consistent with the scenario given by the radiation pressure instability theory. Previous phase-resolved and time-resolved spectral fitting of other sources reached a similar conclusion (Belloni et al., 1997; Neilsen et al., 2011; Rawat et al., 2022).
@@
Besides the phase-resolved fitting of the heartbeat spectra, the energy dependence of the PDS properties and the lags between the photons of different energy bands can also be informative in revealing the origin of the heartbeat. We observe a significant hard lag near the heartbeat characteristic frequency of about one second (similar hard lag at $\sim$ mHz has also been observed in other sources, e.g., GX 339–4 (Uttley et al., 2011) and MAXI J1820+070 (Uttley & Malzac, 2023)). This timescale is several orders of magnitude longer than those expected for reverberation. Instead, a possible explanation for such a large lag is the viscous propagation of mass accretion fluctuations in the disk (Uttley et al., 2011; Lyubarskii, 1997).
@@
Moreover, a positive correlation between fractional rms of the heartbeat variability and energy is observed. This trend is also observed by Yang et al. (2022) to reach an energy of 100 keV. The authors conclude that these Insight-HXMT data constrain the origin of the heartbeat to the corona region. However, our phase-resolved fitting results do not show strong correlation between the count rate and the coronal parameters, which is also the case in the time-resolved fitting of the heartbeat state of GRS 1915+105 in Rawat et al. (2022). Instead, we suggest that disk photons have a temperature modulation, seeding variability which is magnified by the coronal Compton scattering. This could plausibly account for the positive correlation between the fractional rms and energy. A similar scenario has also been proposed in Uttley & Malzac (2024) and supported by analytical calculations.
@@
From the phase-resolved fitting results combined with the radiation pressure instability scenario, we can phenomenologically explain the changes of the inner disk temperature before, during, and after the heartbeat. Before the thermal viscous cycle of the heartbeat begins, the disk accretion process should be on one of the stable branches and the temperature at each radius is locally stable; at the heartbeat cycles onset, the temperature is seen to oscillate over the heartbeat pattern. Because of this, the aggregate spectrum of the heartbeat state blends a range of different thermal components (i.e., a range of disk temperatures). This results in a marked change in the temperature obtained for fitting the time-averaged observation compared to the adjacent data. This is apparent from both the 2021 and 2023 heartbeats states (see Appendix B).
@@ 6.3 The production of wind
A number of BHXRBs have been found to show absorption lines in their spectra which are associated with wind features. Examples include GRS 1915+015 (Zoghbi et al., 2016; Miller et al., 2016; Liu et al., 2022), IGR J17091–-3624 (Wang et al., 2024), MAXI J1803–298 (Zhang et al., 2024), as well as the subject of this paper, 4U 1630–47 (Pahari et al., 2018; Trueba et al., 2019).
There are mainly three candidates to explain the driving mechanism for the wind: the radiation-, thermal-, or magnetic-driving (see Sect. 10 in Done et al., 2007, and references therein, for a general review). In the radiation-driven model, as the luminosity approaches the Eddington-limit, the radiation pressure reduces the effective gravity (by a factor of about $1-L/L_{\mathrm{Edd}}$, where $L_{\mathrm{Edd}}$ is the Eddington luminosity of the source), thus launching the outflow. In the thermal-driven model (Begelman et al., 1983), the central illumination heats the disk to a high temperature, and the heating expands to larger radius due to the pressure gradient. At large enough radii where the thermal energy exceeds the binding energy, a wind will be produced from the outer disk. For the magnetically-driven model (Balbus, 2005a), while magnetohydrodynamic (MHD) simulations shows clear presence of the wind (Balbus, 2005b), a quantitative study of the relationship between the wind and magnetic field is uncertain, e.g., owing to the unknown magnetic-field topology.
@@ 7 Conclusions
    We fit all the spectra with the disk-corona model tbfeo $\times$ thcomp $\otimes$ diskbb and find some spectra showing broadened iron line reflection features in the HIMS and SIMS, and some showing disk wind absorption features in the SIMS and soft state. We fit the reflection spectra with relxillCp and find a stable and untruncated disk in the intermediate states; For the disk wind absorption features, we fit them with zxipcf and find a stable, highly ionized wind with high column density across different outbursts.
@@
    Through the analysis of the light curves and power density spectra, we find two observations in 2021 and 2023 with heartbeat features. The light curves of these observations oscillate quasi-periodically at a frequency around 0.05 Hz and have a heartbeat-like profile. Through phase-resolved fitting of the spectra, we find the flux to be correlated with disk parameters: the inner disk temperature, inner disk radius and mass accretion rate, while no strong correlation with coronal parameters is found. This fast disk oscillation supports the inner disk radiation pressure instability as the driving mechanism of the heartbeats. Moreover, a significant hard lag on the time scale of a second is observed near the characteristic frequency of the heartbeat, which can be explained by the viscous propagation of mass accretion fluctuations in the disk. Combining the phase-resolved spectral analysis and lag analysis, the scenario where the disk-originated oscillation is then magnified by the corona scattering is a possible explanation for the positive relationship between the fractional rms and energy.
@@
Acknowledgments – N.F. thanks Honghui Liu, Zuobin Zhang, Yujia Song, Josephine Wong and Niek Bollemeijer for useful discussions. The work of N.F. and C.B. was supported by the National Natural Science Foundation of China (NSFC), Grant No. 11973019, 12250610185, and 12261131497, and the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400. O.K. acknowledges NICER GO funding 80NSSC23K1660.
@@ Appendix B Heartbeat spectra
Figure 19: The upper panel shows the the spectra of the heartbeat state in 2021 (left panel) and 2023 (right panel), along with spectra from observations taken before and after the heartbeat state. The lower panels show the residuals of the fits of these spectra with the continuum model tbfeo $\times$ thcomp $\otimes$ diskbb.

### v2

@@ 4.2 Relativistic thermal component
Figure 15: As in Fig. 14 but for fits with the spectral model tbfeo $\times$ thcomp $\otimes$ kerrd. Here, the mass accretion rate is shown in the upper panel and the inner disk radius ($R_{\mathrm{in}}$) is shown in the lower panel. The value of $R_{\mathrm{in}}$ in the unit of $R_{\mathrm{g}}$ is measured by setting and distance of the source at 10 kpc and its mass at $10~{}M_{\odot}$ in the fitting. However, the physical meaning of the value should be taken with caution because the source is located near the Galactic plane and has a high column density, therefore limiting a robust constraint on the distance and mass of the source (Seifina et al., 2014).
@@ 5.1 Energy-resolved PDS
We further explore the timing properties of the heartbeat state, first computing the PDS for photons in different energy bands. We fit each with a Lorentzian model, which is commonly used to fit PDS including those with QPOs (Bachetti & Huppenkothen, 2022). For the PDS in the heartbeat state, the variability can be modeled with three Lorentzian functions. We set a zero-centered Lorentzian for the broad band noise, one peaked at the heartbeat characteristic frequency, and a third to fit the noise at higher frequencies. From the central frequency and the full width half maximum (FWHM) of the Lorenzian function corresponding to the heartbeat peak, the quality factor of the 2021 heartbeat oscillation is $2.1^{+0.5}_{-0.9}$, and that of the 2023 heartbeat is $2.6^{+0.9}_{-1.6}$.
@@ 5.2 Energy-resolved time lag and coherence
Analysis of the lags between the photons of different energies is a powerful tool to study the origin of variability in data (Wang et al., 2021; Bollemeijer et al., 2024; Ma et al., 2021), including heartbeats (Yang et al., 2022). In addition, coherence measures the degree of linear correlation between two time series (Bachetti & Huppenkothen, 2022), enabling us to assess the strength of the correlation between the oscillations of light curves of different energy bands. Here we choose the reference energy band to be 2–3 keV. We use stingray to calculate the frequency dependence of the lag and coherence of the chosen energy bands (3–4 keV, 4–7 keV, 7–12 keV) versus the reference energy band in the heartbeat states in 2021 and 2023, and the results are shown in Fig. 17. From the coherence, we can see that the light curves of different energy bands are relatively highly linearly correlated (coherence $\gtrsim 0.8$) near the characteristic frequency of the heartbeats annotated by the yellow region. A hard lag of $\sim 1$ s between 4–12 keV photons and 2–3 keV photons is observed near the heartbeat frequency. Although a few other frequency bins show evidence of similar-amplitude lags (e.g., at $\sim$ 3 mHz in the 2021 and 2023 heartbeat, at $\sim$ 0.105 Hz in the 2023 heartbeat), their coherence is relatively low. The non-linear relationship between signals of different energy bands makes the lag less straightforward to interpret, therefore we do not include these bins into our discussion.
Figure 17: The energy dependence of lag (upper panels) and coherence (lower panels) of the heartbeat states in 2021 (left) and 2023 (right). The reference energy band is 2–3 keV, and the comparison bands are 3–4 keV, 4–7 keV, and 7–12 keV. The yellow region highlights the frequency range of the heartbeat oscillations (from the PDS fits). Near the characteristic frequency of the heartbeats annotated by the yellow region, a hard lag of $\sim 1~{}\mathrm{s}$ between 4–12 keV photons and 2–3 keV photons is observed.
@@ 6 Discussion
6.1 Stable inner disk radius in the soft state
@@
While at lower temperatures the disk emission can generally be well described by a multi-temperature disk blackbody spectrum, non-thermal effects such as the Compton scattering in the disk atmosphere may produce deviations from the multi-temperature disk blackbody spectrum, especially when the inner disk temperature is high. This is normally taken into account by introducing a color correction factor $f_{\mathrm{col}}$ (also called spectral hardening factor) (Shimura & Takahara, 1995; Kubota et al., 1998). Moreover, it has been found in previous studies that the color correction factor has strong dependence on the disk temperature (Kubota & Makishima, 2004; Davis et al., 2006), which can be approximately described by $f_{\mathrm{col}}\propto T_{\mathrm{in}}^{1/4}$. Based on this scaling relation and the equation $norm_{\mathrm{diskbb}}=(R_{\mathrm{in}}/(f_{\mathrm{col}}^{2}D_{10}))^{2}\cos{i}$ (Kubota et al., 1998), we overlay the constant $R_{\mathrm{in}}$ contour in Fig. 12, showing that $R_{\mathrm{in}}$ is essentially constant when the system is in the soft state. The $R_{\rm{in}}$ of the brown points that belong to the 2019 outburst in the bottom-left corner of Fig. 12 looks slightly lower than the $R_{\rm{in}}$ in the soft state in other years of outbursts. Similarly, a slightly lower value of $R_{\rm{in}}$ at the beginning of the 2021 outburst is also found from reflection modeling (see the first two points in Fig. 11). A smaller $R_{\rm{in}}$ compared to the ISCO is not consistent with the standard picture. However, there are a few uncertainties that may be responsible for such a result: $i)$ diskbb is a Newtonian model and can only provide a very rough estimate of the location of the inner edge of the accretion disk, $ii)$ the faintness of the source and the limited quality of these spectra.
Similar results on the inner disk radius have also been found in Swift J1727–1613 (König et al. 2025, in preparation). We note a long-history of evidence from the thermal disk emission for soft states exhibiting a stable accretion-disk inner radius (e.g., detailed analysis of several decades of observations of LMC X–3; see Steiner et al. (2010), and references therein).
@@ 6.2 Heartbeat mechanism
Inner disk radiation instability is a popular candidate mechanism for producing the heartbeat oscillation (Lightman & Eardley, 1974; Done et al., 2007; Neilsen et al., 2012). In that scenario, the dominance of the radiation pressure in the inner disk region and the lack of balancing mechanism give rise to a thermal-viscous instability cycle resulting in the heartbeat oscillations in the light curve (Done et al., 2007; Neilsen et al., 2011).
@@
To gain a better understanding of different physics processes behind this cycle behavior, it is useful to refer to the “S-curve” showing the relationship between the mass and temperature at a certain disk radius (e.g., see Fig. 7 in Done et al., 2007). The curve has three branches: a lower viscous slow rise (where the heating is balanced by radiative cooling and the temperature increases slowly with mass), an upper slow viscous fall (where the heating is balanced by advective cooling and the temperature decreases slowly with mass) and the middle thermal viscous cycle (where the accretion process switches quickly between the two stable branches) (Wang et al., 2024). This cycled oscillation causes a density wave to propagate along the disk and results in the fast changes in the mass accretion rate, inner disk radius, and temperature. The fast variability of disk parameters in our phase-resolved fitting results (Sect. 4) is consistent with the scenario given by the radiation pressure instability theory. Previous phase-resolved and time-resolved spectral fitting of other sources reached a similar conclusion (Belloni et al., 1997; Neilsen et al., 2011; Rawat et al., 2022).
@@
Besides the phase-resolved fitting of the heartbeat spectra, the energy dependence of the PDS properties and the lags between the photons of different energy bands can also be informative in revealing the origin of the heartbeat. We observe a hard lag near the heartbeat characteristic frequency of about one second (similar hard lag at $\sim$ mHz has also been observed in other sources, e.g., GX 339–4 (Uttley et al., 2011) and MAXI J1820+070 (Uttley & Malzac, 2024)). This timescale is several orders of magnitude longer than those expected for reverberation. Instead, a possible explanation for such a large lag is the viscous propagation of mass accretion fluctuations in the disk (Uttley et al., 2011; Lyubarskii, 1997). The broad peaks of the coherence plots may indicate that in a broader frequency range near the heartbeat frequency, photons of different energies are causally related. In the heartbeat state, with the disappearance and refilling of the inner edge of the disk, the characteristic frequency of the viscous propagation lag may slightly change during this process. The frequency range of the coherence peak, corresponding to a time difference of several seconds, is comparable to the viscous time differences for the changing inner disk radius (Belloni et al., 1997).
@@
Moreover, a positive correlation between fractional rms of the heartbeat variability and energy is observed. This trend is also observed by Yang et al. (2022) to reach an energy of 100 keV. The authors conclude that their Insight-HXMT data constrain the origin of the heartbeat to the corona region. However, our phase-resolved fitting results do not show strong correlation between the count rate and the coronal parameters, which is also the case in the time-resolved fitting of the heartbeat state of GRS 1915+105 in Rawat et al. (2022). Instead, we suggest that disk photons have a temperature modulation, seeding variability which is magnified by the coronal Compton scattering. This could plausibly account for the positive correlation between the fractional rms and energy. A similar scenario has also been proposed in Uttley & Malzac (2024) and supported by analytical calculations.
@@
From the phase-resolved fitting results combined with the radiation pressure instability scenario, we can phenomenologically explain the changes of the inner disk temperature before, during, and after the heartbeat. Before the thermal viscous cycle of the heartbeat begins, the disk accretion process should be on one of the stable branches and the temperature at each radius is locally stable; at the heartbeat cycles onset, the temperature is seen to oscillate over the heartbeat pattern. Because of this, the aggregate spectrum of the heartbeat state blends a range of different thermal components (i.e., a range of disk temperatures). This results in a marked change in the temperature obtained for fitting the time-averaged observation compared to the adjacent data. This is apparent from both the 2021 and 2023 heartbeat states (see Appendix B).
In the above discussion, different mechanisms are combined into a more complete scenario of the heartbeat: at a certain strength of mass accretion, the inner disk radiation pressure instability triggers the heartbeat-like oscillations of disk properties. This variation then propagates from the inner part to the outer part of the disk, and extends to higher energy photons when disk photons are scattered by the corona. This scenario highlights the close connection among different parts of the disk-corona system in contributing to the spectral and timing variability seen in the data. Similar discussions about the disk-corona interaction during the heartbeat state of GRS 1915+015 can be found in Yan et al. (2018), and a more general review about that in different systems can be found in Swank (2001).
The precise origin of the different timescales and spectral properties seen between different heartbeat states in the same source is an interesting question but outside of the scope of our work. See Belloni et al. (1997) for a nice discussion about the relationship between the heartbeat period and the radius of the disappearing region in the disk.
@@ 6.3 The production of wind
A number of BHXRBs have been found to show absorption lines in their spectra which are associated with wind features. Examples include GRS 1915+015 (Zoghbi et al., 2016; Miller et al., 2016; Liu et al., 2022), IGR J17091–-3624 (Wang et al., 2024), MAXI J1803–298 (Zhang et al., 2024), as well as the subject of this paper, 4U 1630–47 (Pahari et al., 2018; Trueba et al., 2019). There are mainly three candidates to explain the driving mechanism for the wind: radiation-, thermal-, or magnetic-driving mechanism (see Sect. 10 in Done et al., 2007, and references therein, for a general review and more detailed explanation for different mechanisms).
@@
Based on the averaged wind parameters from the XSTAR fits, the outflowing velocity of the wind is $\sim 0.003~{}c$. Using the spherical approximation, the launching radius of the wind is $\sim 10^{10}~{}\rm{cm}$, corresponding to $\sim 10^{3-4}~{}R_{\rm{g}}$ for a $10~{}M_{\odot}$ black hole. The mass loss rate by the wind is $\sim 10^{18}~{}\rm{g/s}$, corresponding to $\sim 10\%$ Eddington accretion rate for a $10~{}M_{\odot}$ black hole. Therefore, the mass loss rate by the wind is comparable to the mass accretion rate of the source. These estimations of the wind properties are consistent with other studies of the source (Kubota et al., 2007) and comparable to studies of other sources in order of magnitude, for example, GRS 1915+015 (Zoghbi et al., 2016) and MAXI J1803–298 (Zhang et al., 2024). The values of the launching radius and velocity of the wind further support the thermal/magnetic origin of the wind, instead of being radiation-driven.
@@ 7 Conclusions
    We fit all the spectra with the disk-corona model tbfeo $\times$ thcomp $\otimes$ diskbb and find some spectra showing broadened iron line reflection features in the HIMS and SIMS, and some showing disk wind absorption features in the SIMS and soft state. We fit the reflection spectra with relxillCp and find a stable and untruncated disk in the intermediate states; For the disk wind absorption features, we fit them with XSTAR and find a stable, highly ionized wind with high column density across different outbursts.
@@
    Through the analysis of the light curves and power density spectra, we find two observations in 2021 and 2023 with heartbeat features. The light curves of these observations oscillate quasi-periodically at a frequency around 0.05 Hz and have a heartbeat-like profile. Through phase-resolved fitting of the spectra, we find the flux to be correlated with disk parameters: the inner disk temperature, inner disk radius and mass accretion rate, while no strong correlation with coronal parameters is found. This fast disk oscillation supports the inner disk radiation pressure instability as the driving mechanism of the heartbeats. Moreover, a hard lag on the time scale of a second is observed near the characteristic frequency of the heartbeat, which can be explained by the viscous propagation of mass accretion fluctuations in the disk. Combining the phase-resolved spectral analysis and lag analysis, the scenario where the disk-originated oscillation is then magnified by the corona scattering is a possible explanation for the positive relationship between the fractional rms and energy.
@@
Acknowledgments – The authors would like to thank the anonymous referee for valuable suggestions that improved the manuscript. N.F. thanks Honghui Liu, Zuobin Zhang, Yujia Song, Josephine Wong and Niek Bollemeijer for useful discussions. The work of N.F. and C.B. was supported by the National Natural Science Foundation of China (NSFC), Grant No. 11973019, 12250610185, and 12261131497, and the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400. J.F.S and O.K. gratefully acknowledge support from NICER Guest Observer grants 80NSSC23K1096 and 80NSSC24K0535. Y.Z. acknowledges support from the Dutch Research Council (NWO) Rubicon Fellowship, file no. 019.231EN.021. O.K. acknowledges NICER GO funding 80NSSC23K1660.
@@ Appendix B Heartbeat spectra
Figure 19: The upper panel shows the the spectra of the heartbeat state in 2021 (left panel) and 2023 (right panel), along with spectra from observations taken before and after the heartbeat state. The lower panels show the residuals of the fits of these spectra with the model tbfeo $\times$ thcomp $\otimes$ diskbb.